OpenStudy (anonymous):

Quick question because I'm in a bit of rush. k+p=-5 k-p=7 What is k and p?

Mehek (mehek14):

\(k=-5 - p\\ -5 -p-p=7\\-5-2p=7\\5+7=12\\12/-2=-6\\p=-6\\k-6=-5\\6+-5=1\\k=1\)

OpenStudy (anonymous):

Thank you.